9m+n=33
m-4n=16
please help me solve by the elimination method.

Answer by stanbon(75887)

9m+n=33
m-4n=16
--------------
Multiply thru the 1st equation by 4.
36m + 4n = 4*33
m - 4n = 16
------------------
Add and solve for "m":
37m = 4*33+16
m = 4
----------
Substitute for "m" in m-4n=16 and solve for "n":
4 - 4n = 16
-4n = 12
n = -3
============
Cheers,
Stan H.
============